# Focus: Efficient Keyframe Selection for Long Video Understanding

### v1

@@ 3.2 Performance Analysis
Qwen2-VL-7B 32 7B 55.6 57.6
Qwen2-VL-7B w/ Ours 32 7B 62.3 $\uparrow 6.7$ 59.7 $\uparrow 2.1$
@@
This difference helps explain why Focus achieves greater performance gains on LongVideoBench: our method assumes that frame-level relevance scores are i.i.d., a common setting in multi-armed bandit formulations. This assumption neglects temporal dependencies between video segments. Consequently, retrieval-based methods for keyframe selection typically require regularization (Tang et al., 2025; Yu et al., 2024) to promote diversity and ensure coverage.
If temporal dependencies between segments (arms) are taken into account, the problem setting shifts toward Lipschitz or metric bandits (Kleinberg et al., 2008; Bubeck et al., 2011), and contextual bandits (Chu et al., 2011; Agarwal et al., 2014). We leave such extensions to future work.
@@ 3.3 Comparison with State-of-the-Art
  • AKS (Tang et al., 2025): A recent method that adaptively balances frame relevance and temporal coverage. It is considered the current state-of-the-art and also incorporates pre-filtering via downsampling to 1 frame per second (Tang et al., 2025).
@@ 3.4 Efficiency-Accuracy Trade-off
We observe that choice of $\alpha$ has a significant impact on the efficiency while remain stable on the performance. When $\alpha=0.1$, Focus requires around 1.1% of the frames BLIP forward passes while only 3.5 GPU hours. When $\alpha=0.5$, Focus requires around 2.5% of the frames BLIP forward passes while only 9.2 GPU hours. Exhaustively exploiting all arms would require 9.3 GPU hours, while the performance gain compared to $\alpha=0.25$ is negligible.
@@ A.1.1 Multimodal Large Language Models (MLLMs) for Video Understanding
Recent MLLMs extend large language models with visual encoders, encoding images or frames into visual tokens that are fused with text to support open-ended video understanding. Most follow an encode-project-fuse pipeline with instruction tuning, as exemplified by the LLaVA family, Video-LLaVA/Video-LLaMA/Video-ChatGPT, and LLaMA-Vid/VideoChat (Liu et al., 2023; Lin et al., 2023; Zhang et al., 2023; Maaz et al., 2024; Li et al., 2024c; b). Progress has largely come from scaling data/backbones and strengthening cross-modal alignment (MiniCPM-V, InternVL/InternVL2, Qwen2-VL; data-centric and modality-binding advances via ShareGPT4Video and LanguageBind) (Yao et al., 2024; Chen et al., 2024e; d; c; Wang et al., 2024a; Chen et al., 2024a; Zhu et al., 2024), together with architectural refinements that unify multi-granularity visual inputs and tighten temporal adapters, and that improve projector efficiency or curricula (LLaVA-OneVision, LLaVA-NeXT/LLaVA-NeXT-Video, Aria, PLLaVA, Kangaroo) (Li et al., 2025; Liu et al., 2024a; Zhang et al., 2024c; Li et al., 2024a; Xu et al., 2024; Liu et al., 2024b). Finally, several models explicitly target extended context and hierarchical summarization for long-form understanding (LongVILA, LongVA, LongVLM, LongVU) (Chen et al., 2024b; Zhang et al., 2024b; Weng et al., 2024; Shen et al., 2024).
However, this tokenization-first paradigm encounters token explosion on long videos, where dense sampling yields prohibitive sequences. Recent efforts reduce the budget by compressing or restructuring tokens: MovieChat (Song et al., 2024) compacts frames into sparse memory, Video-XL-2 (Qin et al., 2025) synthesizes condensed tokens, and VideoStreaming (Qian et al., 2024) processes streams incrementally to cap tokens. Planning/tool-augmented agents (e.g., VideoAgent (Wang et al., 2024b)) curb perception via selective analysis, while hierarchical controllers (VideoTree (Wang et al., 2025)) and scaling recipes (VideoLLaMA 3 (Zhang et al., 2025a)) aid long-horizon reasoning. Beyond compression, ViLAMP (Cheng et al., 2025) uses mixed-precision tokenization to emphasize differential frames/patches and allocate capacity adaptively; long-context instruction-tuning such as Long-VITA (Shen et al., 2025) complements these strategies for long videos.
@@ A.1.2 Vision-Language Pretrained Models
Cross-modal vision-language pretraining spans two-stream fusion, single-stream fusion, dual-encoder contrastive learning, and encoder-decoder hybrids. Two-stream models such as ViLBERT (Lu et al., 2019) and LXMERT (Tan & Bansal, 2019) encode vision and text separately and fuse via cross-attention, while single-stream counterparts—VisualBERT (Li et al., 2019), VL-BERT (Su et al., 2020), UNITER (Chen et al., 2020)—concatenate region features with text in a unified Transformer using MLM and alignment losses. Large-scale dual encoders like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) learn contrastive embeddings for zero-shot transfer, with FILIP (Yao et al., 2022) improving fine-grained patch-token alignment. Hybrid objectives combine contrastive and generative training (Li et al., 2021; Yu et al., 2022; Wang et al., 2022a; Chen et al., 2023) unify captioning and VQA. The BLIP family integrates vision encoders with language modeling—BLIP (Li et al., 2022) and BLIP-2 (Li et al., 2023) (via a lightweight Q-Former)—while Flamingo (Alayrac et al., 2022) and PaLM-E (Driess et al., 2023) inject visual inputs into large LMs for few-shot multimodal reasoning.
Extending to video, early pretraining models learned joint spatio-temporal-language representations with lightweight fusion and sparse sampling. VideoBERT (Sun et al., 2019) pairs frame sequences with transcripts in a BERT-style objective for retrieval and script generation, while HERO (Li et al., 2020) and ClipBERT (Lei et al., 2021) improve efficiency via hierarchical encoding and key-frame sampling for video-text retrieval and QA. Building directly on large image-text models, Clip4Clip (Luo et al., 2022) reuses CLIP encoders and matches videos to text via contrastive similarity, and FrozenBiLM (Yang et al., 2022) freezes a bi-directional LM while aligning a video encoder for zero-shot VQA.
@@ A.1.3 Keyframe Selection
Recent training-free methods leverage pretrained vision-language models and lightweight heuristics to pick informative, query-relevant frames. Adaptive Keyframe Sampling (AKS) maximizes prompt-frame similarity while enforcing temporal coverage via a split-and-judge policy (Tang et al., 2025); Q-Frame ranks frames by query-conditioned importance and preserves a few at higher resolution for detail (Zhang et al., 2025b). Text-frame alignment with frozen models further enables plug-and-play selectors (KeyVideoLLM, BOLT) that boost Video-LLM performance without fine-tuning (Liang et al., 2024; Liu et al., 2025). To avoid redundancy and preserve structure under a token budget, Logic-in-Frames performs dynamic, logic-verified search (Guo et al., 2025), while VideoTree builds a hierarchical, query-adaptive frame pyramid that expands salient scenes (Wang et al., 2025).
@@
Instruction-guided approaches train selectors with LLM/MLLM feedback: Frame-Voyager learns to query frame combinations by ranking sets with a pretrained Video-LLM (Yu et al., 2024), and Hu et al. (2025) supervise a lightweight selector using MLLM-derived single-frame relevance and multi-frame complementarity. Classical summarization remains relevant: supervised LSTM-based models (vsLSTM, dppLSTM; hierarchical RNNs) learn importance/diversity from human summaries (Zhang et al., 2016; 2018; Zhao et al., 2017), while unsupervised RL/adversarial methods (DR-DSN, SUM-GAN) optimize diversity-representativeness or realism without labels (Zhou et al., 2018; Mahasseni et al., 2017); however, these are typically task-agnostic and may miss frames critical for query-driven VQA.
@@ A.1.4 Multi-Armed Bandits and Batched Exploration
Multi-armed bandits (MAB) encompass both regret minimization and pure exploration. Regret-oriented methods such as UCB variants and Thompson Sampling establish logarithmic-regret foundations for sequential decision-making (Auer et al., 2002; Lai & Robbins, 1985; Agrawal & Goyal, 2012). Pure exploration instead targets high-confidence identification with minimal samples, formalized as best-arm (and top-$k$) identification (Even-Dar et al., 2006; Bubeck et al., 2009; Kalyanakrishnan & Stone, 2010; Cao et al., 2015). Early elimination schemes (Successive/Median Elimination) provide PAC guarantees (Even-Dar et al., 2006; 2002), while confidence-bound and racing families—LUCB, UCB-E, and near-optimal lil’UCB—sharpen sample complexity and approach known lower bounds (Kalyanakrishnan et al., 2012; Audibert & Bubeck, 2010; Karnin et al., 2013; Jamieson et al., 2014; Kaufmann et al., 2016). Beyond single arms, combinatorial pure exploration (CPE) seeks an optimal subset under structural constraints, combining bandit confidence bounds with combinatorial oracles to search exponentially large spaces efficiently (Chen et al., 2016; Lattimore & Szepesvári, 2020).
Fully sequential adaptivity can be impractical when decisions must be made in few rounds or in parallel. Batched (parallel) bandits address this by operating over a small number of adaptivity rounds, yet retain near-sequential sample efficiency for pure exploration in theory and practice (Perchet et al., 2016; Jun et al., 2016; Gao et al., 2019). Batch-elimination/LUCB-style procedures match sequential complexity up to constants with only a handful of updates (Jun et al., 2016), and lower-bound trade-offs between batches and samples are well understood with matching algorithms (Perchet et al., 2016; Kaufmann et al., 2016; Tuynman & Degenne, 2025). Recent designs such as Tri-BBAI attain asymptotically optimal fixed-confidence BAI with just three batches, underscoring the feasibility of resource-constrained exploration (Jin et al., 2024).
@@ Appendix B Bernstein Confidence Radius
$\left|\hat{\mu}_{a}-\mu_{a}\right|\leq\beta_{a}(n)$
@@
$\mathcal{P}\left[\left|\hat{\mu}_{a}-\mu_{a}\right|\leq\sqrt{\frac{2\hat{\sigma}_{a}^{2}\ln\frac{3}{\delta}}{N_{a}(n)}}+\frac{3\ln\frac{3}{\delta}}{N_{a}(n)}\right]\geq 1-2\delta.$
@@ Appendix C Regret Bound
Appendix D Limitations

### v2

@@ 3.2 Performance Analysis
Qwen2-VL-7B 32 7B 55.6 57.4
Qwen2-VL-7B w/ Ours 32 7B 62.3 $\uparrow 6.7$ 59.7 $\uparrow 2.3$
@@
This difference helps explain why Focus achieves greater performance gains on LongVideoBench: our method assumes that frame-level relevance scores are i.i.d., a common setting in multi-armed bandit formulations. This assumption neglects temporal dependencies between video segments. Consequently, retrieval-based methods for keyframe selection typically require regularization (tang2025adaptive; yu2024frame) to promote diversity and ensure coverage.
If temporal dependencies between segments (arms) are taken into account, the problem setting shifts toward Lipschitz or metric bandits (kleinberg2008multi; bubeck2011x), and contextual bandits (chu2011contextual; agarwalb14). We leave such extensions to future work.
@@ 3.3 Comparison with State-of-the-Art
  • AKS (tang2025adaptive): A recent method that adaptively balances frame relevance and temporal coverage. It is considered the current state-of-the-art and also incorporates pre-filtering via downsampling to 1 frame per second (tang2025adaptive).
@@ 3.4 Efficiency-Accuracy Trade-off
We observe that choice of $\alpha$ has a significant impact on the efficiency while remain stable on the performance. With $\alpha=0.1$, Focus evaluates 1.1% of frames and finishes in 3.5 GPU hours. At $\alpha=0.25$, the fraction rises to 1.6% with a cost of 5.5 GPU hours, yielding 63.5% accuracy. Setting $\alpha=0.5$ achieves the highest accuracy (63.6%) but requires evaluating 2.5% of frames and 9.2 GPU hours—only a negligible gain over $\alpha=0.25$ for a substantially higher cost, indicating diminishing returns from exploring more arms.
@@ A.1.1 Multimodal Large Language Models (MLLMs) for Video Understanding
Recent MLLMs extend large language models with visual encoders, encoding images or frames into visual tokens that are fused with text to support open-ended video understanding. Most follow an encode-project-fuse pipeline with instruction tuning, as exemplified by the LLaVA family, Video-LLaVA/Video-LLaMA/Video-ChatGPT, and LLaMA-Vid/VideoChat (liu2023llava; lin2023video; zhang2023videollama; maaz2023videochatgpt; li2023llamavid; li2024videochat). Progress has largely come from scaling data/backbones and strengthening cross-modal alignment (MiniCPM-V, InternVL/InternVL2, Qwen2-VL; data-centric and modality-binding advances via ShareGPT4Video and LanguageBind) (yao2024minicpm; chen2024internvl; chen2024far; chen2024expanding; wang2024qwen2; chen2024sharegpt4video; zhu2024languagebind), together with architectural refinements that unify multi-granularity visual inputs and tighten temporal adapters, and that improve projector efficiency or curricula (LLaVA-OneVision, LLaVA-NeXT/LLaVA-NeXT-Video, Aria, PLLaVA, Kangaroo) (li2024llava; liu2024llavanext; zhang2024llavanextvideo; li2024aria; xu2024pllava; liu2024kangaroo). Finally, several models explicitly target extended context and hierarchical summarization for long-form understanding (LongVILA, LongVA, LongVLM, LongVU) (longvila; zhang2024longva; weng2024longvlm; shen2024longvu).
However, this tokenization-first paradigm encounters token explosion on long videos, where dense sampling yields prohibitive sequences. Recent efforts reduce the budget by compressing or restructuring tokens: MovieChat (song2024moviechat) compacts frames into sparse memory, Video-XL-2 (qin2025video) synthesizes condensed tokens, and VideoStreaming (qian2024streaming) processes streams incrementally to cap tokens. Planning/tool-augmented agents (e.g., VideoAgent (wang2024videoagent)) curb perception via selective analysis, while hierarchical controllers (VideoTree (wang2025videotree)) and scaling recipes (VideoLLaMA 3 (zhang2025videollama)) aid long-horizon reasoning. Beyond compression, ViLAMP (cheng2025scaling) uses mixed-precision tokenization to emphasize differential frames/patches and allocate capacity adaptively; long-context instruction-tuning such as Long-VITA (shen2025long) complements these strategies for long videos.
@@ A.1.2 Vision-Language Pretrained Models
Cross-modal vision-language pretraining spans two-stream fusion, single-stream fusion, dual-encoder contrastive learning, and encoder-decoder hybrids. Two-stream models such as ViLBERT (lu2019vilbert) and LXMERT (tan2019lxmert) encode vision and text separately and fuse via cross-attention, while single-stream counterparts—VisualBERT (li2019visualbert), VL-BERT (su2020vlbert), UNITER (chen2020uniter)—concatenate region features with text in a unified Transformer using MLM and alignment losses. Large-scale dual encoders like CLIP (radford2021learning) and ALIGN (jia2021scaling) learn contrastive embeddings for zero-shot transfer, with FILIP (yao2022filip) improving fine-grained patch-token alignment. Hybrid objectives combine contrastive and generative training (li2021align; yu2022coca; wang2022ofa; chen2022pali) unify captioning and VQA. The BLIP family integrates vision encoders with language modeling—BLIP (li2022blip) and BLIP-2 (li2023blip2) (via a lightweight Q-Former)—while Flamingo (alayrac2022flamingo) and PaLM-E (driess2023palme) inject visual inputs into large LMs for few-shot multimodal reasoning.
Extending to video, early pretraining models learned joint spatio-temporal-language representations with lightweight fusion and sparse sampling. VideoBERT (sun2019videobert) pairs frame sequences with transcripts in a BERT-style objective for retrieval and script generation, while HERO (li2020hero) and ClipBERT (lei2021clipbert) improve efficiency via hierarchical encoding and key-frame sampling for video-text retrieval and QA. Building directly on large image-text models, Clip4Clip (luo2022clip4clip) reuses CLIP encoders and matches videos to text via contrastive similarity, and FrozenBiLM (yang2022zero) freezes a bi-directional LM while aligning a video encoder for zero-shot VQA.
@@ A.1.3 Keyframe Selection
Recent training-free methods leverage pretrained vision-language models and lightweight heuristics to pick informative, query-relevant frames. Adaptive Keyframe Sampling (AKS) maximizes prompt-frame similarity while enforcing temporal coverage via a split-and-judge policy (tang2025adaptive); Q-Frame ranks frames by query-conditioned importance and preserves a few at higher resolution for detail (zhang2025q). Text-frame alignment with frozen models further enables plug-and-play selectors (KeyVideoLLM, BOLT) that boost Video-LLM performance without fine-tuning (liang2024keyvideollm; liu2025bolt). To avoid redundancy and preserve structure under a token budget, Logic-in-Frames performs dynamic, logic-verified search (guo2025logic), while VideoTree builds a hierarchical, query-adaptive frame pyramid that expands salient scenes (wang2025videotree).
@@
Instruction-guided approaches train selectors with LLM/MLLM feedback: Frame-Voyager learns to query frame combinations by ranking sets with a pretrained Video-LLM (yu2024frame), and hu2025m supervise a lightweight selector using MLLM-derived single-frame relevance and multi-frame complementarity. Classical summarization remains relevant: supervised LSTM-based models (vsLSTM, dppLSTM; hierarchical RNNs) learn importance/diversity from human summaries (zhang2016video; zhang2018recurrent; zhao2017hierarchical), while unsupervised RL/adversarial methods (DR-DSN, SUM-GAN) optimize diversity-representativeness or realism without labels (zhou2018deep; mahasseni2017unsupervised); however, these are typically task-agnostic and may miss frames critical for query-driven VQA.
@@ A.1.4 Multi-Armed Bandits and Batched Exploration
Multi-armed bandits (MAB) encompass both regret minimization and pure exploration. Regret-oriented methods such as UCB variants and Thompson Sampling establish logarithmic-regret foundations for sequential decision-making (auer2002ucb; lai1985asymptotically; agrawal2012analysis). Pure exploration instead targets high-confidence identification with minimal samples, formalized as best-arm (and top-$k$) identification (even2006action; bubeck2009pure; kalyanakrishnan2010efficient; cao2015top). Early elimination schemes (Successive/Median Elimination) provide PAC guarantees (even2006action; even2002pac), while confidence-bound and racing families—LUCB, UCB-E, and near-optimal lil’UCB—sharpen sample complexity and approach known lower bounds (kalyanakrishnan2012pac; audibert2010best; karnin2013almost; jamieson2014lil; kaufmann2016complexity). Beyond single arms, combinatorial pure exploration (CPE) seeks an optimal subset under structural constraints, combining bandit confidence bounds with combinatorial oracles to search exponentially large spaces efficiently (chen2016combinatorial; lattimore2020bandit).
Fully sequential adaptivity can be impractical when decisions must be made in few rounds or in parallel. Batched (parallel) bandits address this by operating over a small number of adaptivity rounds, yet retain near-sequential sample efficiency for pure exploration in theory and practice (perchet2016batched; jun2016top; gao2019batched). Batch-elimination/LUCB-style procedures match sequential complexity up to constants with only a handful of updates (jun2016top), and lower-bound trade-offs between batches and samples are well understood with matching algorithms (perchet2016batched; kaufmann2016complexity; tuynman2025batch). Recent designs such as Tri-BBAI attain asymptotically optimal fixed-confidence BAI with just three batches, underscoring the feasibility of resource-constrained exploration (jin2024optimal).
@@ Appendix B Bernstein Confidence Radius
$\left|\hat{\mu}_{a}(n)-\mu_{a}\right|\leq\beta_{a}(n)$
@@
$\mathcal{P}\left[\left|\hat{\mu}_{a}(n)-\mu_{a}\right|\leq\sqrt{\frac{2\hat{\sigma}_{a}^{2}\ln\frac{3}{\delta}}{N_{a}(n)}}+\frac{3\ln\frac{3}{\delta}}{N_{a}(n)}\right]\geq 1-2\delta.$
@@ Appendix C Regret Bound
Appendix D Visualizations of Failure Cases
Figure 4: Two representative failure modes of LLaVA-Video-7B when using Focus to select keyframes. Yellow stars mark manually annotated frames that are most informative for the query. In the first case, Focus correctly selects these frames, but the MLLM still fails to answer due to its limited ability to reason over the relatively complex chart. In the second case, Focus fails to capture the critical frames during a compact, rapid scene transition: the relevant segment lasts only 1-2 seconds within a 10-minute video, making the keyframes difficult to identify even for human experts.
To provide a more comprehensive understanding of the proposed Focus, we analyze two typical failure patterns of LLaVA-Video-7B when using Focus to select keyframes in Figure 4, which most failure cases fall into.
In the first case, the query asks: “When a pie chart representing the Czech ethnicity appears in the video, with blue occupying the largest portion, red the second, and light green the least, which of the following sentences is displayed on the screen?” Across the entire video, this pie chart appears multiple times and is interleaved with other background content. Consequently, even though Focus correctly selects the most informative frames, the MLLM is confused by the subtle differences between multiple similar pie charts. This failure pattern is mainly attributable to the limited reasoning and perception capabilities of the MLLM itself, rather than to the keyframe selection method.
In the second case, the video is a 10-minute vlog with frequent scene transitions. The query asks: “In a room with green wall tiles, there is a woman with long hair wearing a white dress. In the lower part of the screen near her head, white text appears that says ‘someone started playing drums in the back.’ What change happens to her when she appears in the restroom?” The relevant segment lasts only 1–2 seconds within the 10-minute video, making the keyframes difficult to identify even for human experts. As shown in Figure 4, Focus successfully selects frames where the correct text appears, but still fails to capture the most critical frames. This pattern reveals that, in some intrinsically challenging cases, the adaptive sampling strategy of Focus may risk missing crucial information.
Appendix E Comparison with State-of-the-Art
Model #Frame LLM LongVideoBench Video-MME
Qwen2-VL-7B 32 7B 55.6 57.4
Qwen2-VL-7B w/ AKS 32 7B 57.8 59.7
Qwen2-VL-7B w/ Q-Frame 32 7B 57.4 56.5
Qwen2-VL-7B w/ Ours 32 7B 62.3 $\uparrow 6.7$ 59.7 $\uparrow 2.3$
LLaVA-OV-7B 32 7B 54.8 56.5
LLaVA-OV-7B w/ AKS 32 7B 57.4 57.7
LLaVA-OV-7B w/ Q-Frame 32 7B 54.8 56.8
LLaVA-OV-7B w/ Ours 32 7B 60.7 $\uparrow 5.9$ 58.3 $\uparrow 1.8$
LLaVA-Video-7B 64 7B 58.9 64.4
LLaVA-Video-7B w/ AKS 64 7B 62.1 64.6
LLaVA-Video-7B w/ Q-Frame 64 7B 59.9 64.5
LLaVA-Video-7B w/ Ours 64 7B 63.5 $\uparrow 4.6$ 65.4 $\uparrow 1.0$
Table 5: Video question-answering accuracy (%) of different MLLMs on LongVideoBench and Video-MME. We compare Focus with AKS and Q-Frame on Qwen2-VL, LLaVA-OV, and LLaVA-Video. The suffix “w/ Ours” denotes models using keyframes selected by Focus; likewise, “w/ AKS” and “w/ Q-Frame” indicate using keyframes from the corresponding baselines. #Frame is the number of frames fed into the MLLM, and LLM denotes the language model size.
Here we compare our proposed Focus against state-of-the-art training-free keyframe selection methods on both LongVideoBench and Video-MME. Specifically, we consider two recent approaches based on vision-language similarity:
  • AKS (tang2025adaptive): A plug-and-play adaptive keyframe sampling module that recursively balances query–frame relevance and temporal coverage under a fixed frame budget. By first downsampling the video to 1 frame per second, scoring each frame with a prompt–frame matching model, and then applying a judge-and-split procedure to allocate keyframe slots across segments, AKS maximizes informative coverage and serves as a strong state-of-the-art baseline for long-video QA.
  • Q-Frame (zhang2025q): A training-free, query-aware frame selection and multi-resolution adaptation framework that can be plugged in front of diverse Video-LLMs. It uses a text–image matching network (e.g., CLIP) to compute query–frame similarity scores, samples a compact set of highly relevant frames via stochastic selection, and assigns them heterogeneous resolutions so that crucial frames are preserved at high fidelity under a fixed token budget.
We report the results in Table 5. Across all three backbones and both benchmarks, Focus consistently outperforms both AKS and Q-Frame under the same frame budget. In particular, Focus improves the plain Qwen2-VL-7B, LLaVA-OV-7B, and LLaVA-Video-7B models by 4.6–6.7% on LongVideoBench and up to 2.3% on Video-MME, indicating that our keyframe selection strategy transfers robustly across different MLLMs.
For the two compared baselines, AKS consistently outperforms Q-Frame on both LongVideoBench and Video-MME whenever Q-Frame is evaluated. We attribute this to the more sophisticated and adaptive sampling scheme of AKS, which explicitly balances query–frame relevance and temporal coverage instead of relying solely on similarity scores.
By contrast, Q-Frame behaves more like a token-compression mechanism: it maps a fixed frame budget to a fixed number of visual tokens so that the MLLM can "see" more frames than it is originally designed for. However, the lack of an explicit temporal sampling or search design means that Q-Frame does not actively reason about where informative moments occur in long videos, which limits its performance in the long-form setting.
Appendix F Experiments on More Benchmarks
Model #Frame LLM MLVU VSI-Bench
Qwen2-VL-7B 32 7B 59.7 36.5
Qwen2-VL-7B w/ AKS 32 7B 64.3 36.9
Qwen2-VL-7B w/ Ours 32 7B 67.0 $\uparrow 6.7$ 39.0 $\uparrow 2.5$
LLaVA-Video-7B 64 7B 68.2 41.7
LLaVA-Video-7B w/ AKS 64 7B 71.2 42.2
LLaVA-Video-7B w/ Ours 64 7B 72.7 $\uparrow 4.5$ 42.4 $\uparrow 0.7$
Table 6: Video question-answering accuracy (%) of different MLLMs on MLVU and VSI-Bench. We compare Focus with AKS on Qwen2-VL and LLaVA-Video. The suffix “w/ Ours” denotes models using keyframes selected by Focus; likewise, “w/ AKS” indicates using keyframes from the corresponding baselines. #Frame is the number of frames fed into the MLLM, and LLM denotes the language model size.
To further investigate the generalization ability of Focus beyond long-form QA benchmarks, we conduct experiments on two additional datasets:
  • MLVU (zhou2025mlvu): A comprehensive multi-task long-video understanding benchmark constructed from 1,730 long videos (3 minutes to 2 hours) spanning movies, surveillance, egocentric recordings, cartoons, and game videos. It defines nine evaluation tasks that jointly probe both global and local reasoning abilities of MLLMs, and reveals substantial performance degradation as video length grows.
  • VSI-Bench (yang2025thinking): A video-based visual–spatial intelligence benchmark built from 288 egocentric indoor videos (ScanNet, ScanNet++, ARKitScenes) with over 5,000 question–answer pairs. It focuses on 3D spatial understanding and memory from first-person streams, evaluating MLLMs on tasks such as spatial layout reasoning, navigation, and distance estimation.
We summarize the results in Table 6. On MLVU, our method improves Qwen2-VL-7B from 59.7% to 67.0% (+7.3%) and LLaVA-Video-7B from 68.2% to 72.7% (+4.5%), while also outperforming AKS by +2.7% and +1.5% points, respectively. On VSI-Bench, which emphasizes fine-grained spatial reasoning over relatively short egocentric clips, our method still yields consistent gains: for Qwen2-VL-7B, accuracy increases from 36.5% to 39.0% (+2.5%), and for LLaVA-Video-7B from 41.7% to 42.4% (+0.7%), respectively. These results indicate that our temporal search mechanism generalizes well across different backbones and tasks, with particularly pronounced benefits on long and heterogeneous videos.
At the same time, the improvements on VSI-Bench are understandably smaller than on long-video benchmarks. When videos are short and informative content is more uniformly distributed, uniform sampling already captures many salient frames, leaving less headroom for sophisticated temporal search. We explicitly regard this as a limitation and a promising direction for future work on spatially-aware frame selection in low-redundancy settings.
Appendix G Ablation Studies
G.1 Two-Stage Exploration-Exploitation
One of the core designs of Focus is the two-stage exploration-exploitation procedure. To better understand the contribution of each stage, we introduce two variants of Focus:
  • Focus-C: This variant only performs the coarse exploration stage to identify promising temporal arms. In the final keyframe selection step, it randomly samples frames from all frames within the selected arms without any further refinement.
  • Focus-F: This variant only performs the fine-grained exploration stage by uniformly sampling frames over the whole video and interpolating the rewards via nearest-neighbor assignment. The final keyframes are then drawn directly from the resulting video-level sampling distribution, without the arm-level pre-selection.
Uniform FOCUS-C FOCUS-F FOCUS
Qwen2-VL 55.6 61.7 61.5 62.3
LLaVA-OV 54.8 58.4 57.7 60.7
LLaVA-Video 58.9 62.3 62.5 63.5
Table 7: Ablation of the two-stage exploration-exploitation procedure on LongVideoBench. Uniform denotes naive uniform frame sampling. Focus-C uses only the coarse exploration stage to select promising temporal arms, and then randomly samples frames within them. Focus-F uses only the fine-grained exploration stage over the entire video. The full Focus combines both stages and consistently achieves the best performance across all MLLMs, indicating that coarse arm selection and fine-grained refinement are complementary.
We conduct experiments on LongVideoBench with Qwen2-VL-7B, LLaVA-OV-7B, and LLaVA-Video-7B, and summarize the ablation results in Table 7. Both Focus-C and Focus-F provide substantial improvements over uniform sampling across all three backbones, demonstrating that coarse arm selection and fine-grained exploration are each effective on their own. The full two-stage variant further yields the best performance in all cases, achieving an additional gain of up to 2.3% over the single-stage variants, which confirms that coarse localization of promising regions and subsequent fine-grained exploitation are complementary rather than interchangeable.
G.2 Bernstein confidence radius
Compared with the classical UCB algorithm, the Bernstein confidence radius is more robust to high-variance rewards. To better understand its contribution, we introduce a variant of Focus that relies on the empirical mean without a variance-aware exploration bonus when selecting top-relevance frames:
  • Focus-M: This variant uses the empirical mean reward to rank arms and select top-relevance frames, instead of the Bernstein confidence radius.
Uniform FOCUS-M FOCUS
Qwen2-VL 55.6 61.7 62.3
LLaVA-OV 54.8 58.1 60.7
LLaVA-Video 58.9 63.0 63.5
Table 8: Ablation of the Bernstein confidence radius on LongVideoBench. Uniform denotes naive uniform frame sampling. FOCUS-M uses the empirical mean to rank arms and select top-relevance frames. The full FOCUS leverages the Bernstein confidence radius to form variance-aware upper confidence bounds.
We summarize the results in Table 8. The empirical-mean variant (FOCUS-M) already yields large gains over uniform sampling across all three backbones, showing that even a simple bandit-style selection is beneficial. However, the full method with the Bernstein confidence radius consistently achieves the best performance, providing up to 2.6% improvement over uniform and up to 2.6% improvement over the base models. This confirms that a variance-aware confidence radius is more effective than the empirical mean alone for selecting top-relevance frames, as it encourages additional exploration of high-uncertainty clips, especially when a clip contains diverse or rapidly changing scenes.
G.3 Effect of Clip Length
In the formulation of Focus, each video is partitioned into fixed-length clips that serve as bandit arms. The clip length $l$ is a crucial hyper-parameter that controls the granularity of exploration and exploitation. To better understand its effect, we conduct experiments on LongVideoBench with LLaVA-Video-7B and summarize the results in Table 9.
Uniform 8s 16s 32s
ACC 58.9 63.7 63.5 62.3
GPU hours 8.1 5.5 4.1
Table 9: Ablation of the clip length $l$ on LongVideoBench with LLaVA-Video-7B. Uniform denotes naive uniform frame sampling (thus no additional GPU hours for keyframe selection are reported). For Focus, we vary the clip length from 8s to 32s and report both QA accuracy and the GPU hours required for keyframe selection. Note that the GPU hours are measured on a single NVIDIA H100 (80GB) GPU.
As shown in Table 9, all clip-length settings of Focus significantly outperform uniform sampling (58.9% vs. 62.3–63.7%), indicating that our bandit-based selection is robust to the choice of $l$ over a reasonably wide range. Shorter clips (e.g., 8s) provide slightly better accuracy by enabling more fine-grained exploration, but they also incur higher computational cost, while longer clips (e.g., 32s) reduce GPU hours at the price of a modest performance drop. In practice, we find $l=16$ seconds to offer a good trade-off between accuracy and efficiency.
G.4 Effect of Vision-Language Encoder
Our method can be seamlessly integrated with different vision-language encoders to estimate frame-query relevance scores. In the main experiments, we adopt BLIP to align with our primary baseline AKS for a fair comparison, and also because prior work has shown BLIP to be a robust and effective choice for frame-level relevance estimation. To provide a more comprehensive evaluation, we further conduct experiments with three encoders: CLIP (radford2021learning), SigLIP (zhai2023sigmoid), and BLIP (li2022blip).
Uniform CLIP SigLIP BLIP
ACC 58.9 60.2 60.9 63.5
Table 10: Ablation of the vision-language encoder on LongVideoBench with LLaVA-Video-7B. Uniform denotes naive uniform frame sampling. For our method, we instantiate the frame-query scoring module with CLIP, SigLIP, and BLIP.
As summarized in Table 10, all three encoders yield clear improvements over uniform sampling, confirming that our bandit-based selection is compatible with different vision-language backbones. Among them, BLIP achieves the strongest performance, while CLIP and SigLIP still provide 1.3% and 2.0% gains, respectively. These results suggest that our framework is robust to the choice of encoder, but can further benefit from stronger frame-query relevance models, and that future advances in vision-language pretraining are likely to directly translate into better keyframe selection performance.
Appendix H Limitations